# TOB-SVD: Total-Order Broadcast with Single-Vote Decisions in the Sleepy Model

### v2

@@ 5.1 Graded Agreement with $k=2$ grades
Finally, similarly to Integrity, Uniqueness follows from the fact that outputting $\Lambda$ with any grade requires a majority of unique log senders to have sent a log that extends $\Lambda$, without counting logs received from equivocators. This ensures that the sets of senders of logs which a validator counts in support of conflicting logs do not intersect, and so that a validator cannot see a majority for two conflicting logs. We only go through the argument for grade 00, because Uniqueness for grade $1$ is already implied by Consistency. Note first that there is a natural injection of $V^{2\Delta,i}_{\Lambda}$ into $S^{2\Delta,i}$, since at most one log per validator is considered in the former, due to removing logs from equivocators. Moreover, note that $V^{2\Delta,i}_{\Lambda^{\prime}}$ and $V^{2\Delta,i}_{\Lambda}$ are disjoint for conflicting $\Lambda$ and $\Lambda^{\prime}$, so $|V^{2\Delta,i}_{\Lambda^{\prime}}|+|V^{2\Delta,i}_{\Lambda}|\leq|S^{2\Delta,i}|$. Therefore, $|V^{2\Delta,i}_{\Lambda}|>|S^{2\Delta,i}|/2$ implies $|V^{2\Delta,i}_{\Lambda^{\prime}}|\leq|S^{2\Delta,i}|/2$ for any conflicting $\Lambda^{\prime}$. ∎
@@ 5.2 Graded Agreement with $k=3$ grades
We now analyze the Graded Agreement protocol with $k=3$ grades.
This protocol lasts $5\Delta$ time instead of $3\Delta$, and requires $|H_{0,0,5\Delta}|>|B_{5\Delta}|$, similarly to the previous GA, but with $5\Delta$ instead of $3\Delta$ for $T_{b}$. Therefore, it in particular works in the $(5\Delta,0,1)$-sleepy model.
During the output phase for grade $1$, honest validators that are awake at time $4\Delta$ only participate if they were also active at time $2\Delta$. In the output phase for grade $2$, validators active at time $5\Delta$ will only participate if they were awake at time $\Delta$. This is because outputting a log with either grade $1$ or $2$ requires having previously stored supporting logs for them, as per the time-shifted quorum technique.
@@
The proof for all the properties of the Graded Agreement with $k=3$ grades are similar to those for the Graded Agreement with $k=2$ grades (Theorem 1). For this reason, we only discuss the Graded Delivery property, which is where the key idea of the protocol is utilized, i.e., a nested application of the time-shifted quorum technique. Let $v_{i}$, $v_{j}$, and $v_{k}$ be three honest validators participating in the output phase for grade 00, grade $1$, and grade $2$, respectively. Time $[2\Delta,4\Delta]$ in this GA functions exactly like time $[\Delta,3\Delta]$ in the Graded Agreement with $k=2$ grades: at first, logs $V^{2\Delta}$ are stored for later use, then comes the output phase for grade 00, where a log $\Lambda$ is output with grade 00 if $|V^{3\Delta}_{\Lambda}|>|S^{3\Delta}|/2$, and finally comes the output phase for grade $1$, where $\Lambda$ is output with grade $1$ if $|V^{2\Delta}_{\Lambda}\cap V^{4\Delta}_{\Lambda}|>|S^{4\Delta}|/2$. In other words, for grade 00 and grade $1$ we have a first application of the time-shifted quorum technique. For the same reasons as in the Graded Agreement with $k=2$ grades, we then have that $|V^{2\Delta,j}_{\Lambda}\cap V^{4\Delta,j}_{\Lambda}|\leq|V^{3\Delta,i}_{\Lambda}|$ and $|S^{3\Delta,i}|\leq|S^{4\Delta,j}|$, which guarantees the Graded Delivery property from grade $1$ to grade 00. This application of the time-shifted quorum technique is nested inside another such application, which guarantees the Graded Delivery property from grade $2$ to grade $1$. Firstly, the participation level for grade $2$ outputs, i.e., $S^{5\Delta}$, is determined time $\Delta$ after that for grade $1$ outputs, $S^{4\Delta}$, which ensures the correct inclusion, i.e., $S^{4\Delta,j}\subseteq S^{5\Delta,k}$. Conversely, $V^{\Delta}_{\Lambda}$, the initial supporting logs for grade $2$ outputs, are determined time $\Delta$ before $V^{2\Delta}_{\Lambda}$, those for grade $1$ outputs. Finally, the sets of equivocating senders, whose votes are discarded from the supporting logs, are determined in the same order as the participation level, to ensure that any sender which is considered an equivocator in output phase of grade $1$ is also considered an equivocator in the output phase for grade $2$. Together, these last two points guarantee that $V^{\Delta,k}_{\Lambda}\cap V^{5\Delta,k}_{\Lambda}\subseteq V^{2\Delta,j}_{\Lambda}\cap V^{4\Delta,j}_{\Lambda}$, so that $|V^{\Delta,k}_{\Lambda}\cap V^{5\Delta,k}_{\Lambda}|>|S^{5\Delta,k}|/2$ implies $|V^{2\Delta,j}_{\Lambda}\cap V^{4\Delta,j}_{\Lambda}|\geq|V^{\Delta,k}_{\Lambda}\cap V^{5\Delta,k}_{\Lambda}|>|S^{5\Delta,k}|/2\geq|S^{4\Delta,j}|/2$, i.e., validator $v_{k}$ outputting $(\Lambda,2)$ implies validator $v_{j}$ outputting $(\Lambda,1)$. ∎
6 TOB-SVD
We now present our dynamically available Total-Order Broadcast protocol – TOB-SVD – built upon the Graded Agreement with $k=3$ grades. Our approach takes inspiration from 1/4MMR [8], which also uses a Graded Agreement with $k=3$ grades to reduce the number of invocations of Graded Agreement to one per view.
The protocol, which is presented in Figure 4, proceeds in views of $4\Delta$ time each. We let $t_{v}=4\Delta v$ be the beginning of view $v$. To each view $v$ corresponds a Graded Agreement $GA_{v}$, which runs in the time interval $[t_{v}+\Delta,t_{v}+6\Delta]=[t_{v}+\Delta,t_{v+1}+2\Delta]$, i.e., $GA_{v}$ takes up some of view $v+1$ as well. Moreover, the GA invocations are not perfectly sequential, as $GA_{v}$ and $GA_{v+1}$ overlap during time $[t_{v+1}+\Delta,t_{v+1}+2\Delta]$.
At the beginning of each view $v$ there is a proposal phase corresponding to the output phase for grade 00 of $GA_{v-1}$. The leader of view $v$ proposes a log $\Lambda$ extending its grade 00 output (a candidate), if it has one.
This is followed by a decision phase at time $t_{v}+2\Delta$, where logs proposed (at best) in the previous view ($v-1$) can be decided. In particular, this phase corresponds to the output phase for grade $2$ of $GA_{v-1}$, and such grade $2$ outputs are decided. This is safe, because Graded Delivery of Graded Agreement ensures that all honest validators (among those participating in the output phase for grade $1$) output with grade $1$ a decided log, thus locking on it and therefore inputting to $GA_{v}$ a log extending it. During the decision phase, validators also update $V$ with all $\mathsf{LOG}$ messages received until that point in time for $GA_{v}$ from non-equivocating validators, as part of the ongoing $GA_{v}$.
Finally, no further action is taken at time $t_{v}+3\Delta$ other than what is required by $GA_{v}$ itself, i.e., to store in $V^{2\Delta}$ all $\mathsf{LOG}$ messages received until that point from non-equivocating validators.
Outputs of $GA_{-1}$, which are used in view $v=0$, are all formally defined to be the log containing only the Genesis log. In each view $v\geq 0$, awake validators participate in the GA instances that are ongoing, and in addition behave as specified here whenever they have the required GA outputs to do so. Validators do not perform actions which require outputs they do not have. 1. Propose ($t=t_{v}$): Output phase for grade 00 of $GA_{v-1}$. Propose $\Lambda^{\prime}$ extending $\Lambda$, the highest log output with grade 00 by $GA_{v-1}$ (candidate), accompanied by the VRF value for view $v$. 2. Vote ($t=t_{v}+\Delta$): Output phase for grade $1$ of $GA_{v-1}$. $GA_{v}$ starts. Let $L_{v-1}$ be the highest log output with grade $1$ by $GA_{v-1}$ (lock). After discarding equivocating proposals, input to $GA_{v}$ the proposal with the highest VRF value extending $L_{v-1}$, or $L_{v-1}$ if no such proposal exists. 3. Decide ($t=t_{v}+2\Delta$): Output phase for grade $2$ of $GA_{v-1}$, which ends. Decide the highest log $\Lambda$ output with grade $2$ by $GA_{v-1}$. 4. ($t=t_{v}+3\Delta$): Do nothing (other than what is required by the ongoing $GA_{v}$).
@@ 6 TOB-SVD
In order to prove this theorem, we first ensure that the properties of our Graded Agreement primitive (Figure 2) hold when employing it within the Total-Order Broadcast protocol (Figure 4).
Observe that, due to the assumed stabilization period of $2\Delta$, our Total-Order Broadcast works in the $(5\Delta,2\Delta,1)$-sleepy model, i.e., when $|H_{t-2\Delta,t,t+5\Delta}|>|B_{t+5\Delta}|$ holds for every time $t\geq 0$. On the other hand, the Graded Agreement protocol itself (Figure 2) works in the $(5\Delta,0,1)$-sleepy model, requiring $|H_{t,t,t+5\Delta}|>|B_{t+5\Delta}|$, without any stability assumption ($T_{s}=0$). When invoking it as part of our Total-Order Broadcast, the only change is in the input phase, since honest validators use the outputs of $GA_{v-1}$ to determine the inputs to $GA_{v}$. In particular, validators in $H_{t_{v}+\Delta}$ input something to $GA_{v}$ if and only if they have a lock $L_{v-1}$, such that they output $L_{v-1}$ with grade $1$ in $GA_{v-1}$. For that to be the case, they must have participated in the output phase for grade $1$ of $GA_{v-1}$, which requires them to also have been awake at time $t_{v-1}+3\Delta=t_{v}-\Delta$. The additional stability requirement of $T_{s}=2\Delta$ in the Total-Order Broadcast model takes care of this, ensuring that we only consider those honest validators which are allowed to participate in the input phase of a $GA_{v}$, i.e., validators in $H_{t-2\Delta,t}$, for $t=t_{v}+\Delta$.
@@
Suppose an honest validator $v_{i}$ decides log $\Lambda$ at time $t_{v}+2\Delta$ by outputting $(\Lambda,2)$ in $GA_{v-1}$. By the Graded Delivery property of Graded Agreement, any honest validator participating in the output phase for grade $1$ of $GA_{v-1}$ outputs $(\Lambda,1)$. By Lemma 1, for any $v^{\prime}\geq v-1$, honest validators participating in the output phase for grade $1$ of $GA_{v^{\prime}}$ output $(\Lambda,1)$. Now, suppose that another honest validator $v_{j}$ decides a conflicting $\Lambda^{\prime}$ and, without loss of generality, let us assume that $v_{j}$ does so during view $v^{\prime\prime}\geq v$. Again, by the Graded Delivery property, every honest validator which participates in the output phase of $GA_{v^{\prime\prime}-1}$ outputs $(\Lambda^{\prime},1)$. Since $v^{\prime\prime}-1\geq v-1$ we have shown that any such validator also outputs $(\Lambda,1)$, contradicting the Uniqueness property of Graded Agreement. ∎
@@
Observe that $|H_{t_{v}}\setminus{B_{t_{v}+\Delta}}|\geq|H_{t_{v}-2\Delta,t_{v},t_{v}+5\Delta}|>|B_{t_{v}+5\Delta}|\geq|B_{t_{v}+\Delta}|$. Thus, $2|H_{t_{v}}\setminus{B_{t_{v}+\Delta}}|>|B_{t_{v}+\Delta}|+|H_{t_{v}}\setminus{B_{t_{v}+\Delta}}|=|H_{t_{v}}\cup B_{t_{v}+\Delta}|$. Equivalently, $|H_{t_{v}}\setminus{B_{t_{v}+\Delta}}|>\frac{1}{2}|H_{t_{v}}\cup B_{t_{v}+\Delta}|$. View $v$ has a good leader whenever a validator in $H_{t_{v}}\setminus{B_{t_{v}+\Delta}}$ has the highest VRF value for view $v$ out all validators in $H_{t_{v}}\cup B_{t_{v}+\Delta}$. The adversary is mildly adaptive, so corruptions which happen by time $t_{v}+\Delta$ must have been scheduled by time $t_{v}$. In particular, the adversary has to determine $H_{t_{v}}\cap B_{t_{v}+\Delta}$ before observing any of the VRF values of validators in $H_{t_{v}}$. Therefore, view $v$ has a good leader with probability $\frac{|H_{t_{v}}\setminus{B_{t_{v}+\Delta}}|}{|H_{t_{v}}\cup B_{t_{v}+\Delta}|}>\frac{1}{2}$. ∎
@@
Consider any such honest validator $v_{i}$ and its lock $L_{v-1}$, which $v_{i}$ outputs with grade $1$ in $GA_{v-1}$. Note that by the Validity property and the fact that any log is an extension of $\Lambda_{\mathsf{g}}$, any honest validator participating in the output phase for grade $1$ outputs some log extending $\Lambda_{\mathsf{g}}$. As the leader $v_{\ell}$ is honest and awake at time $t_{v}$, by the Graded Delivery property of Graded Agreement, $v_{\ell}$ outputs $(L_{v-1},0)$ in $GA_{v-1}$ , and, by the Uniqueness property, it does not output any conflicting log with grade 00. This means that the proposal $\Lambda$ made by leader $v_{\ell}$ extends $L_{v-1}$. The proposal is received by validator $v_{i}$ by time $t_{v}+\Delta$, and no other proposal from $v_{\ell}$ is received by $v_{i}$ at that point, because the leader is still honest at time $t_{v}+\Delta$, since $v_{\ell}\not\in B_{t_{v}+\Delta}$ by definition of a good leader. Moreover, no other proposal received by $v_{i}$ at this point has higher VRF value, since a good leader for view $v$ has the highest VRF value among all validators from which a proposal might have been received ($H_{t_{v}}\cup B_{t_{v}+\Delta})$. Therefore, validator $v_{i}$ inputs $\Lambda$ to $GA_{v}$. ∎
Theorem 5.
@@
[…]

### v3

@@ 6.1 Graded Agreement with $k=2$ grades
Finally, similarly to Integrity, Uniqueness follows from the fact that outputting $\Lambda$ with any grade requires a majority of unique log senders to have sent a log that extends $\Lambda$, without counting logs received from equivocators. This ensures that the sets of senders of logs which a validator counts in support of conflicting logs do not intersect, and so that a validator cannot see a majority for two conflicting logs. We only go through the argument for grade 0, because Uniqueness for grade $1$ is already implied by Consistency. Note first that there is a natural injection of $V^{2\Delta,i}_{\Lambda}$ into $S^{2\Delta,i}$, since at most one log per validator is considered in the former, due to removing logs from equivocators. Moreover, note that $V^{2\Delta,i}_{\Lambda^{\prime}}$ and $V^{2\Delta,i}_{\Lambda}$ are disjoint for conflicting $\Lambda$ and $\Lambda^{\prime}$, so $|V^{2\Delta,i}_{\Lambda^{\prime}}|+|V^{2\Delta,i}_{\Lambda}|\leq|S^{2\Delta,i}|$. Therefore, $|V^{2\Delta,i}_{\Lambda}|>|S^{2\Delta,i}|/2$ implies $|V^{2\Delta,i}_{\Lambda^{\prime}}|\leq|S^{2\Delta,i}|/2$ for any conflicting $\Lambda^{\prime}$. ∎
@@ 6.2 Graded Agreement with $k=3$ grades
We now analyze the Graded Agreement protocol with $k=3$ grades. Recall that this protocol works in the $(5\Delta,0,\frac{1}{2})$-sleepy model.
During the output phase for grade $1$, honest validators that are awake at time $4\Delta$ only participate if they were also awake at time $2\Delta$. In the output phase for grade $2$, validators awake at time $5\Delta$ will only participate if they were awake at time $\Delta$. This is because outputting a log with either grade $1$ or $2$ requires having previously stored supporting logs for them, as per the time-shifted quorum technique.
@@
The proof for all the properties of the Graded Agreement with $k=3$ grades are similar to those for the Graded Agreement with $k=2$ grades (Theorem 1). For this reason, we only discuss the Graded Delivery property, which is where the key idea of the protocol is utilized, i.e., a nested application of the time-shifted quorum technique. Let $v_{i}$, $v_{j}$, and $v_{k}$ be three honest validators participating in the output phase for grade 0, grade $1$, and grade $2$, respectively. Time $2\Delta,4\Delta]$ in this GA functions exactly like time $\Delta,3\Delta]$ in the Graded Agreement with $k=2$ grades: at first, logs $V^{2\Delta}$ are stored for later use, then comes the output phase for grade 0, where a log $\Lambda$ is output with grade 0 if $|V^{3\Delta}_{\Lambda}|>|S^{3\Delta}|/2$, and finally comes the output phase for grade $1$, where $\Lambda$ is output with grade $1$ if $|V^{2\Delta}_{\Lambda}\cap V^{4\Delta}_{\Lambda}|>|S^{4\Delta}|/2$. In other words, for grade 0 and grade $1$ we have a first application of the time-shifted quorum technique. For the same reasons as in the Graded Agreement with $k=2$ grades, we then have that $|V^{2\Delta,j}_{\Lambda}\cap V^{4\Delta,j}_{\Lambda}|\leq|V^{3\Delta,i}_{\Lambda}|$ and $|S^{3\Delta,i}|\leq|S^{4\Delta,j}|$, which guarantees the Graded Delivery property from grade $1$ to grade 0. This application of the time-shifted quorum technique is nested inside another such application, which guarantees the Graded Delivery property from grade $2$ to grade $1$. Firstly, the participation level for grade $2$ outputs, i.e., $S^{5\Delta}$, is determined time $\Delta$ after that for grade $1$ outputs, $S^{4\Delta}$, which ensures the correct inclusion, i.e., $S^{4\Delta,j}\subseteq S^{5\Delta,k}$. Conversely, $V^{\Delta}_{\Lambda}$, the initial supporting logs for grade $2$ outputs, are determined time $\Delta$ before $V^{2\Delta}_{\Lambda}$, those for grade $1$ outputs. Finally, the sets of equivocating senders, whose votes are discarded from the supporting logs, are determined in the same order as the participation level, to ensure that any sender which is considered an equivocator in output phase of grade $1$ is also considered an equivocator in the output phase for grade $2$. Together, these last two points guarantee that $V^{\Delta,k}_{\Lambda}\cap V^{5\Delta,k}_{\Lambda}\subseteq V^{2\Delta,j}_{\Lambda}\cap V^{4\Delta,j}_{\Lambda}$, so that $|V^{\Delta,k}_{\Lambda}\cap V^{5\Delta,k}_{\Lambda}|>|S^{5\Delta,k}|/2$ implies $|V^{2\Delta,j}_{\Lambda}\cap V^{4\Delta,j}_{\Lambda}|\geq|V^{\Delta,k}_{\Lambda}\cap V^{5\Delta,k}_{\Lambda}|>|S^{5\Delta,k}|/2\geq|S^{4\Delta,j}|/2$, i.e., validator $v_{k}$ outputting $(\Lambda,2)$ implies validator $v_{j}$ outputting $(\Lambda,1)$. ∎
6.3 TOB-SVD
We conclude by analyzing our main result, TOB-SVD.
@@ 6.3 TOB-SVD
In order to prove this theorem, we first ensure that the properties of our Graded Agreement primitive (Figure 2) hold when employing it within the TOB-SVD protocol (Figure 4).
Observe that, due to the assumed stabilization period of $2\Delta$, our Total-Order Broadcast works in the $(5\Delta,2\Delta,\frac{1}{2})$-sleepy model, i.e., when $|B_{t+5\Delta}|<|H_{t-2\Delta,t}\cup B_{t+5\Delta}|/2$ holds for every time $t\geq 0$. On the other hand, the Graded Agreement protocol itself (Figure 2) works in the $(5\Delta,0,\frac{1}{2})$-sleepy model, requiring $|B_{t+5\Delta}|<|H_{t,t}\cup B_{t+5\Delta}|/2$, without any stability assumption ($T_{s}=0$). When invoking it as part of our Total-Order Broadcast, the only change is in the input phase, since honest validators use the outputs of $GA_{v-1}$ to determine the inputs to $GA_{v}$. In particular, validators in $H_{t_{v}+\Delta}$ input something to $GA_{v}$ if and only if they have a lock $L_{v-1}$, such that they output $L_{v-1}$ with grade $1$ in $GA_{v-1}$. For that to be the case, they must have participated in the output phase for grade $1$ of $GA_{v-1}$, which requires them to also have been awake at time $t_{v-1}+3\Delta=t_{v}-\Delta$. The additional stability requirement of $T_{s}=2\Delta$ in the Total-Order Broadcast model takes care of this, ensuring that we only consider those honest validators which are allowed to participate in the input phase of a $GA_{v}$, i.e., validators in $H_{t-2\Delta,t}$, for $t=t_{v}+\Delta$.
@@
Suppose an honest validator $v_{i}$ decides log $\Lambda$ at time $t_{v}+2\Delta$ by outputting $(\Lambda,2)$ in $GA_{v-1}$. By the Graded Delivery property, any honest validator participating in the output phase for grade $1$ of $GA_{v-1}$ outputs $(\Lambda,1)$. By Lemma 1, for any $v^{\prime}\geq v-1$, honest validators participating in the output phase for grade $1$ of $GA_{v^{\prime}}$ output $(\Lambda,1)$. Now, suppose that another honest validator $v_{j}$ decides a conflicting $\Lambda^{\prime}$ and, without loss of generality, let us assume that $v_{j}$ does so during view $v^{\prime\prime}\geq v$. Again, by the Graded Delivery property, every honest validator which participates in the output phase of $GA_{v^{\prime\prime}-1}$ outputs $(\Lambda^{\prime},1)$. Since $v^{\prime\prime}-1\geq v-1$ we have shown that any such validator also outputs $(\Lambda,1)$, contradicting the Uniqueness property of Graded Agreement. ∎
@@
Observe that, due to Condition (1), $\rho=\frac{1}{2}$ and $B_{t_{v}}\subseteq B_{t_{v}+5\Delta}$, $|H_{t_{v}}\setminus{B_{t_{v}+\Delta}}|\geq|H_{t_{v}}\setminus{B_{t_{v}+5\Delta}}|=|H_{t_{v}}\cup{B_{t_{v}+5\Delta}}|-|{B_{t_{v}+5\Delta}}|>|H_{t_{v}}\cup{B_{t_{v}+5\Delta}}|/2\geq|H_{t_{v}}\cup{B_{t_{v}+\Delta}}|/2$. View $v$ has a good leader whenever a validator in $H_{t_{v}}\setminus{B_{t_{v}+\Delta}}$ has the highest VRF value for view $v$ out all validators in $H_{t_{v}}\cup B_{t_{v}+\Delta}$. The adversary is mildly adaptive, so corruptions which happen by time $t_{v}+\Delta$ must have been scheduled by time $t_{v}$. In particular, the adversary has to determine $H_{t_{v}}\cap B_{t_{v}+\Delta}$ before observing any of the VRF values of validators in $H_{t_{v}}$. Therefore, view $v$ has a good leader with probability $\frac{|H_{t_{v}}\setminus{B_{t_{v}+\Delta}}|}{|H_{t_{v}}\cup B_{t_{v}+\Delta}|}>\frac{1}{2}$. ∎
@@
Consider any such honest validator $v_{i}$ and its lock $L_{v-1}$, which $v_{i}$ outputs with grade $1$ in $GA_{v-1}$. Note that by the Validity property and the fact that any log is an extension of $\Lambda_{\mathsf{g}}$, any honest validator participating in the output phase for grade $1$ outputs some log extending $\Lambda_{\mathsf{g}}$. As the leader $v_{\ell}$ is honest and awake at time $t_{v}$, by the Graded Delivery property of Graded Agreement, $v_{\ell}$ outputs $(L_{v-1},0)$ in $GA_{v-1}$ , and, by the Uniqueness property, it does not output any conflicting log with grade 0. This means that the proposal $\Lambda$ made by leader $v_{\ell}$ extends $L_{v-1}$. The proposal is received by validator $v_{i}$ by time $t_{v}+\Delta$, and no other proposal from $v_{\ell}$ is received by $v_{i}$ at that point, because the leader is still honest at time $t_{v}+\Delta$, since $v_{\ell}\not\in B_{t_{v}+\Delta}$ by definition of a good leader. Moreover, no other proposal received by $v_{i}$ at this point has higher VRF value, since a good leader for view $v$ has the highest VRF value among all validators from which a proposal might have been received by ($H_{t_{v}}\cup B_{t_{v}+\Delta})$. Therefore, validator $v_{i}$ inputs $\Lambda$ to $GA_{v}$. ∎
Lemma 4.
@@
Take a valid transaction $\mathit{tx}$ in the pool of transactions. By Lemma 2 and our assumption on transaction batching by honest validators, there exists a view $v$ with an honest leader which proposes a log $\Lambda$ that includes $\mathit{tx}$. As per Lemma 4, $\mathit{tx}$ is included in the log decided by any honest validator that after $t_{v+1}-2\Delta$ is eventually awake for at least $8\Delta$. ∎